# Enhanced Visual Question Answering: A Comparative Analysis and Textual Feature Extraction Via Convolutions

### v1

@@ Abstract
Visual Question Answering (VQA) has emerged as a highly engaging field in recent years, attracting increasing research efforts aiming to enhance VQA accuracy through the deployment of advanced models such as Transformers. Despite this growing interest, there has been limited exploration into the comparative analysis and impact of textual modalities within VQA, particularly in terms of model complexity and its effect on performance. In this work, we conduct a comprehensive comparison between complex textual models that leverage long dependency mechanisms and simpler models focusing on local textual features within a well-established VQA framework. Our findings reveal that employing complex textual encoders is not invariably the optimal approach for the VQA-v2 dataset. Motivated by this insight, we introduce an improved model, ConvGRU, which incorporates convolutional layers to enhance the representation of question text. Tested on the VQA-v2 dataset, ConvGRU achieves better performance without substantially increasing parameter complexity.
Keywords Visual Question Answering  $\cdot$ Textual Features  $\cdot$ N-Grams  $\cdot$ Convolution
@@ 1 Introduction
Visual Question Answering (VQA) has emerged as an increasingly interesting area of research at the intersection of computer vision and natural language processing. The objective of VQA is to answer questions about a given image, requiring models capable of understanding both visual and textual modalities. One particular task in the VQA domain that merits additional attention is the counting task. This specific task requires the model to quantitatively identify the number of objects in a given image. Despite its apparent simplicity, this task poses significant challenges, as it demands precise identification, localization, and counting of objects, all while avoiding potential pitfalls like multiple counting or not acknowledging occluded items.
Advanced sequential models, such as Transformers, have seen staggering success in various tasks and have been widely adopted. These models, known for their ability to handle long-range dependencies and complex patterns, have become somewhat the gold standard in the field. However, their effectiveness in all types of tasks, particularly in those that are relatively simple and exhibit closely similar semantics, like VQA tasks, is not always guaranteed.
@@
In this paper, we pose the question: Are complex sequential models the most suitable approach for handling textual modality in VQA tasks, especially on the original VQA-v2 dataset? To investigate this, we conducted rigorous experiments applying both complex models like Transformer Encoder and attention-based models, as well as simpler structures such as RNNs and CNNs on the VQA-v2 dataset. We focused our analysis on different text feature extractions and their impact on model accuracy. We discovered that simpler models, specifically those adept at capturing local interdependencies within the text, could provide notable improvements over some complex models. Furthermore, the incorporation of convolutional layers into simpler structures like Gated Recurrent Units (GRU), forming what we term as ConvGRU, proved quite effective.
The subsequent sections will detail our methods, experimental setup, evaluation metrics, and results. We also conduct an analysis of textual length of the VQA-v2 dataset, and demonstrate the effectiveness of local feature extraction through several qualitative cases. Our findings suggest that embracing simplicity can often lead to improved efficiency and accuracy, offering insights that challenge common practices within the field of VQA tasks.
Figure 1: Overview of the VQA architecture, where the innovative component is the ConvGRU module highlighted in green. The question text is first converted into embeddings, which are then processed through convolutional layers of different kernel sizes (2-gram and 3-gram features) to capture local semantic information. These features are concatenated and fed into a GRU unit to integrate temporal information. The image features are extracted through ResNet-101, followed by using bounding boxes and spatial attention to focus on key areas of the image.Through a fusion module, the image and text information are combined, and finally, the classifier outputs the answer. Aside from the design of the ConvGRU, the yellow modules utilize the same model configurations and structures as [1].
@@ 2 Related Work
Visual Question Answering. Visual Question Answering (VQA) has rapidly evolved since [2] first introduced the concept of free-form and open-ended VQA tasks, accompanied by a benchmark dataset and evaluation metrics. [3] introduces complementary image-question pairs to mitigate the imbalance caused by language priors. Early VQA research primarily utilized Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) for image and question encoding, respectively [2][4][5][6]. [4] exemplifies this approach through a two-branch neural network combining a CNN-based image encoder with an LSTM question encoder. The emergence of Transformer models has then catalyzed the development of cross-modal architectures in the VQA domain. Notably, models like ViLBERT [7], VisualBERT [8], and CLIP [9] have led the transition towards unified frameworks. These models have been instrumental in recent VQA research efforts [10] [11] [12] [13], showcasing remarkable results on VQA tasks by effectively bridging visual and textual modalities.
@@
Spatial Attention in VQA. Spatial attention mechanism is adopted in order to improve the performance of VQA and mimic human behavior, i.e., looking at certain parts of the regions [14][15][16][17]. Stacked attention networks (SANs) [14] and Spatial Memory Network(SMem)[15] similarly utilize two-stage attention framework to produce better glimpses through multiple reasoning. To generate a query-aware context representation without early summarization, [16] proposes a multi-stage hierarchical bi-directional attention network. [17] introduces a unique neural network to simulate the attention with a multivariate distribution while taking into account spatial interactions between regions.
@@
Text Representations in VQA. The task of Visual Question Answering (VQA) necessitates effective textual representation to comprehend questions that guide the visual understanding. Traditional approaches have predominantly employed Recurrent Neural Networks (RNNs), including Gated Recurrent Units (GRU) and Long Short-Term Memory (LSTM) networks [2] [4] [18]. Their ability to model the temporal dependencies inherent in natural language questions makes them robust baselines for text representation in VQA tasks. Beyond RNNs, the Text Convolutional Neural Network (TextCNN) [19] emerges as an alternative that leverages convolutional operations to extract local features, offering a different perspective on capturing textual semantics [5]. The advent of the Transformer [20] has introduced a shift towards attention-based structures for text representation. The Transformer Encoder, with its self-attention mechanism, allows for the direct modeling of relationships between all words in a question, irrespective of their positional distances. Building on the success of Transformers, BERT (Bidirectional Encoder Representations from Transformers) [21] has emerged as another milestone in text representation adept at grasping the nuanced context of words in a sentence from both directions. This progression has culminated in Transformers and BERT becoming the dominant models in recent VQA works [10] [22].
Image Captioning. Image captioning is a multi-modal task similar to VQA. It aims at giving a description in natural language based on the visual content of the input image. The most common approaches are based on neural encoders (of images, usually CNNs), and decoders (of captions, usually LSTM), and attention mechanism is adopted to focus on specific regions of images, such as stochastic-attention and deterministic-attention models in [23]. [24] devises a meshed-memory Transformer via a mesh connectivity between encoder layers and decoder layers. Different modifications are applied to improve pre-training Transformer models, such as X-linear attention block in [25]. LayoutLMv2 in [26] integrate text, layout and images via alignment and matching subtasks. Regarding object detectors, cross-modal retrieval module and image conditioning module are utilized in [27].
@@ 3.2 Method Overview
Our methodology seeks to improve VQA task outcomes by focusing on enhanced text processing techniques, leveraging the foundational work of [1]. We introduce a model that primarily innovates in textual feature extraction through a novel Convolutional GRU framework, while also efficiently integrating pre-existing advanced methods for visual analysis and a counting component to address both visual data and numerical queries effectively.
At the heart of our proposed model, as depicted in Figure 1, is the integration of a Convolutional GRU framework for improved extraction of textual features from questions. For image analysis, we employ a pre-trained ResNet-101 model for extracting features, complemented by a stacked attention mechanism and a counting module from the base model. This combination offers a comprehensive solution for the VQA challenge.
@@ 3.3.1 Image Encoder
We employ a pre-trained ResNet-101 model [28] for image feature extraction, leveraging its architecture optimized for depth and computational efficiency through "Bottleneck" blocks. To capture image features directly, we exclude the model’s final pooling and fully connected layers. Table 1 provides an overview of ResNet-101’s structure and the arrangement of its Bottleneck blocks.
Table 1: Detailed Structure of ResNet-101 Used for Image Feature Extraction. The table outlines the configuration of ResNet-101’s layers, including convolutional layers, max pooling, and the "Bottleneck" blocks arrangement.
@@ 3.3.2 Stacked Attention Mechanism
This process produces an attention map $x_{t}$, which is then employed to calculate a focused weighted sum of image features:
@@ 3.3.3 Counting Module
Adopted from [1], the counting module integrates attention weights $a$ and and bounding boxes $b$ as input to enable precise object counting. The process unfolds through several key operations, streamlined for clarity:
@@ 3.3.4 Fusion and Classifier Module
For classification, a multimodal approach is utilized, concatenating visual and question representations, along with the counting module’s output. This concatenated result undergoes $ReLU$ activation and is further refined through batch normalization before proceeding to a fully-connected layer designed for a 3000-category classification task.
@@ 3.4 Convolutional GRU for Text Encoding
The extraction of textual features in Visual Question Answering tasks has traditionally relied on Recurrent Neural Networks (RNNs), particularly favoring Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU) for their efficacy. GRUs, with their streamlined architecture, have been preferred for VQA due to the generally straightforward semantic nature of questions and their lower computational cost. Our main contribution lies in enhancing the GRU architecture by incorporating convolutionally extracted n-gram features as inputs instead of direct word embeddings. This innovation is driven by the insight that specific keywords within questions play a crucial role in determining the answers, mirroring the function of attention mechanisms in processing visual inputs.
@@
Considering the concise nature of VQA questions, our analysis reveals that self-attention mechanisms might lead to overfitting, a limitation observed in our comparative analysis with RCNN+SA-GRU models (Section 4.3). We propose a one-dimensional convolutional approach as a more effective alternative for feature extraction. By employing kernels of sizes 2 and 3, we capture bi-gram and tri-gram features, respectively, which are then processed through distinct padding strategies-asymmetrical for bi-grams to prioritize the first word of the sequence, and symmetrical for tri-grams to preserve context evenly. Similar to the observations in [29] regarding the shift problem encountered with even-numbered convolutions in image processing, we identify a parallel issue in the realm of one-dimensional text convolutions. However, to avoid the potential complexity and overfitting that might arise from directly addressing this shift problem, we adopt a strategic approach by adding a unit of padding at the beginning of sequences. This single padding enhances our model’s ability to accurately identify the initial words of a question, which are pivotal in defining its category-for instance, "what" typically signals a query about an object, while "what color" suggests a question about color. The utility of this asymmetric head padding technique is also validated by the results in Section 4.3.
@@
Figure 2: Schematic representation of the textual modality feature extraction model. Word Embeddings are first applied through a Dropout layer and Tanh activation, and subsequently processed through asymmetric head-padding and symmetric padding. Convolutional layers (Conv2 and Conv3) capture multi-scale local features, and a residual unit incorporating shortcut connections is applied to enhance the feature set, ensuring even questions with subtle n-gram features retain robust initial representations. These enhanced features are then concatenated, forming a rich representation that feeds into a gated recurrent unit (GRU). The GRU then effectively extracts sequential semantic information from these local features. The inset details the GRU’s internal gating mechanisms that regulate information flow, with gates for resetting $r_{i}$ and updating $z_{i}$, and the candidate activation $\tilde{h}_{i}$, which, together with the current state $h_{i-1}$, influences the output state $h_{i}$.
Figure 2 outlines our Convolutional GRU (ConvGRU) model’s architecture, highlighting the integration of convolutional n-gram feature extraction with GRU processing.
@@
Given a sequence of question words $q$, each word is first embedded into a high-dimensional space, yielding the matrix $E\in\mathbb{R}^{batch\_size\times seq\_length\times embed\_size}$. Following a dropout and Tanh activation, the matrix is prepared for convolution:
@@
With this schema, our Convolutiona GRU improves upon traditional methods by adeptly capturing and emphasizing pivotal textual features, thus providing a robust foundation for more accurate responses.
@@ 4.1 Dataset and Evaluation Metric
where $A$ is the human answer set including 10 answers, and $Agree(\cdot)$ is the number of human answers same with the predicted one. An answer is considered correct if at least 3 humans give the same answer.
To account for the imbalance of answers in the dataset, our study predominantly employs the Balanced Pairs Accuracy [30] metric. This optimizes evaluation fairness by forming "complementary image pairs" and a model needs to correctly answer on both to get a score.
@@ 4.2 Experimental Settings
Training Details. All experiments described in this study were conducted on a single RTX 3090 GPU with 24GB of memory. For the purpose of maintaining consistency and fairness across all tests, we utilized uniform parameters when evaluating a variety of Text Feature Extraction Models. Specifically, the following settings were applied to each experiment:
@@
    $lr_{i+1}=lr_{i}\times 0.5^{(1/lr\_halflife)}$ (16)
    Here, $lr_{i}$ represents the learning rate at the current epoch, and $lr_{i+1}$ denotes the learning rate for the next epoch. The term $lr\_halflife$ is set to 50,000, indicating that the learning rate is halved every 50,000 epochs.
@@
Compared Methods. Image features are computed from ResNet101 mentioned in Section 3.3.1. Different Text models are adopted, including basic one-layer Gated Recurrent Unit (GRU), Text Convolutional Neural Network (TextCNN), two-layer Long Short-Term Memory Network(2Layer-LSTM), Bidirectional GRU (BiGRU), Bidirectional LSTM (BiLSTM), Self-Attention GRU (SA-GRU), Multi-Head Self-Attention GRU (MHA-GRU), Transformer Encoder (TE) and Transformer Encoder with GRU (TE-GRU). Specific settings are as follows:

### v2

# Enhanced Textual Feature Extraction for Visual Question Answering: A Simple Convolutional Approach

@@ Abstract
Visual Question Answering (VQA) has emerged as a highly engaging field in recent years, with increasing research focused on enhancing VQA accuracy through advanced models such as Transformers. Despite this growing interest, limited work has examined the comparative effectiveness of textual encoders in VQA, particularly considering model complexity and computational efficiency. In this work, we conduct a comprehensive comparison between complex textual models that leverage long-range dependencies and simpler models focusing on local textual features within a well-established VQA framework. Our findings reveal that employing complex textual encoders is not invariably the optimal approach for the VQA-v2 dataset. Motivated by this insight, we propose ConvGRU, a model that incorporates convolutional layers to improve text feature representation without substantially increasing model complexity. Tested on the VQA-v2 dataset, ConvGRU demonstrates a modest yet consistent improvement over baselines for question types such as Number and Count, which highlights the potential of lightweight architectures for VQA tasks, especially when computational resources are limited.
Keywords Visual Question Answering  $\cdot$ Textual Feature Extraction  $\cdot$ N-Grams  $\cdot$ Convolution
@@ 1 Introduction
Visual Question Answering (VQA) [1] has emerged as an increasingly interesting area of research at the intersection of computer vision and natural language processing. The objective of VQA is to answer questions about a given image, requiring models capable of understanding both visual and textual modalities. One particular task in the VQA domain that merits additional attention is the counting task. This specific task requires the model to quantitatively identify the number of objects in a given image. Despite its apparent simplicity, this task poses significant challenges, as it demands precise identification, localization, and counting of objects, all while avoiding potential pitfalls like multiple counting or not acknowledging occluded items.
Advanced sequential models, such as Transformers, have seen staggering success in various tasks and have been widely adopted. However, these models are not always the best choice for every task, particularly in settings where balancing model performance with computational cost is essential. As highlighted in previous studies that evaluate models under specific conditions [2], carefully considering model selection and exploring simple yet effective improvements is also crucial for optimizing task performance.
@@
In this paper, we pose the question: Are complex sequential models the most suitable approach for handling textual modality in VQA tasks, especially on the original VQA-v2 dataset? To investigate this problem, we conduct rigorous experiments applying both complex models like Transformer Encoder and attention-based models, as well as simpler structures such as RNNs and CNNs on the VQA-v2 dataset. We focus our analysis on different text feature extractions and their impact on model accuracy. We discover that simpler models, specifically those good at capturing local interdependencies within the text, can provide improvements over some complex models. Furthermore, the incorporation of convolutional layers into simpler structures like Gated Recurrent Units (GRU), forming what we term as ConvGRU, proves quite feasible.
Our main contributions are threefold: (1) We present ConvGRU, a GRU-based text encoder enhanced by utilizing convolutional layers. Experimental results demonstrate that leveraging local textual features with ConvGRU can marginally improve VQA accuracy without increasing computational costs. (2) We conduct an extensive comparison of various VQA text encoders, revealing that complex models such as Transformer Encoders may underperform compared to simpler architectures better at extracting local textual features.
Figure 1: Overview of the VQA architecture highlighting the ConvGRU module. Questions are embedded and processed through convolutional layers to capture local semantics, then integrated via a GRU unit. Image features are extracted using ResNet-101 with spatial attention focusing on key areas. The fusion module combines image and text information, and the classifier outputs the answer. Counting, attention and classifier modules follow the configurations of.
@@ 2 Related Work
Visual Question Answering. Visual Question Answering (VQA) has rapidly evolved since [1] first introduced the concept of free-form and open-ended VQA tasks, accompanied by a benchmark dataset and evaluation metrics. [3] introduces complementary image-question pairs to mitigate the imbalance caused by language priors. Early VQA research primarily utilized Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) for image and question encoding, respectively [1, 4, 5, 6]. [4] exemplifies this approach through a two-branch neural network combining a CNN-based image encoder with an LSTM question encoder. The emergence of Transformer models has then catalyzed the development of cross-modal architectures in the VQA domain. Notably, models like ViLBERT [7], VisualBERT [8], and CLIP [9] have led the transition towards unified frameworks. These models have been instrumental in recent VQA research efforts [10, 11, 12, 13], showcasing remarkable results on VQA tasks.
@@
Spatial Attention in VQA. Spatial attention mechanism is adopted in order to improve the performance of VQA and mimic human behavior, i.e., looking at certain parts of the regions [14, 15, 16]. Stacked attention networks (SAN) [14] and Spatial Memory Network (SMem) [15] similarly utilize two-stage attention framework to produce better glimpses through multiple reasoning. To generate a query-aware context representation without early summarization, [16] proposes a multi-stage hierarchical bi-directional attention network.
@@
Text Representations in VQA. The task of Visual Question Answering (VQA) necessitates effective textual representation to comprehend questions that guide the visual understanding. Traditional approaches have predominantly employed Recurrent Neural Networks (RNNs), including Gated Recurrent Units (GRU) and Long Short-Term Memory (LSTM) networks [1, 4, 17]. Their ability to model the temporal dependencies inherent in natural language questions makes them robust baselines for text representation in VQA tasks. Beyond RNNs, the Text Convolutional Neural Network (TextCNN) [18] emerges as an alternative that leverages convolutional operations to extract local features, offering a different perspective on capturing textual semantics [5]. The advent of the Transformer [19] has introduced a shift towards attention-based structures for text representation. The Transformer Encoder, with its self-attention mechanism, allows for the direct modeling of relationships between all words in a question, irrespective of their positional distances. Building on the success of Transformers, BERT (Bidirectional Encoder Representations from Transformers) [20] has emerged as another milestone in text representation adept at grasping the nuanced context of words in a sentence from both directions. This progression has culminated in Transformers and BERT becoming the dominant models in recent VQA works [10, 21].
@@ 3.2 Method Overview
Our methodology seeks to improve VQA task outcomes by focusing on enhanced text processing techniques, leveraging the foundational work of [22]. We introduce a model that primarily innovates in textual feature extraction through a novel Convolutional GRU framework, while also efficiently integrating pre-existing advanced methods for image feature extraction and a counting component to address both visual data and numerical queries effectively.
At the heart of our proposed model, as depicted in Figure 1, is the integration of a Convolutional GRU framework for improved extraction of textual features from questions. For image modality, we employ a pre-trained R-CNN model for extracting features, complemented by a stacked attention mechanism and a counting module from the base model. This combination offers a comprehensive solution for the VQA challenge.
@@ 3.3.1 Image Encoder
We employ a pre-trained R-CNN model with a ResNet-101 backbone [23] for image feature extraction, leveraging its architecture optimized for depth and computational efficiency through bottleneck blocks. Table 1 provides an overview of ResNet-101 structure and the arrangement of its bottleneck blocks.
Table 1: Detailed Structure of ResNet-101 Used for Image Feature Extraction. The table outlines the configuration of layers in ResNet-101, including convolutional layers, max pooling, and the bottleneck blocks.
@@ 3.3.2 Stacked Attention Mechanism
The attention map $x_{t}$ is then employed to calculate a focused weighted sum of image features:
@@ 3.3.3 Counting Module
Adopted from [22], the counting module integrates attention weights $a$ and bounding boxes $b$ as input to enable precise object counting. The process unfolds through several key operations, streamlined for clarity:
@@ 3.3.4 Fusion and Classifier Module
For classification, a multi-modal approach is utilized, concatenating visual and question representations, along with the counting module’s output. This concatenated result undergoes $ReLU$ activation and is further refined through batch normalization before proceeding to a fully-connected layer designed for a 3000-category classification task.
@@ 3.4 Convolutional GRU for Text Encoding
In Visual Question Answering (VQA), textual feature extraction traditionally utilizes Recurrent Neural Networks (RNNs) like Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU) due to their proficiency in modeling sequential data. GRUs are often favored for their computational efficiency, especially given the concise and straightforward nature of questions in VQA tasks. Our main contribution lies in enhancing the GRU architecture by incorporating convolutionally extracted n-gram features as inputs instead of direct word embeddings. This approach is theoretically grounded in the understanding that local n-gram patterns are crucial for capturing semantic meaning in text [24].
@@
We propose a one-dimensional convolutional approach as a more effective alternative for feature extraction. By employing kernels of sizes 2 and 3, we capture bi-gram and tri-gram features, respectively, which are then processed through distinct padding strategies-asymmetrical for bi-grams to prioritize the first word of the sequence, and symmetrical for tri-grams to preserve context evenly. Similar to the observations in [25] regarding the shift problem encountered with even-numbered convolutions in image processing, we identify a parallel issue in the realm of one-dimensional text convolutions. To avoid the potential complexity and overfitting that might arise from directly addressing this shift problem, we choose to simply add a unit of padding at the beginning of sequences to accurately identify the initial words of a question, which are crucial for defining its category. For instance, "what" typically signals a query about an object, while "what color" suggests a question about color. The utility of this asymmetric head padding technique is also validated by the results in Section 4.3.
@@
Figure 2: Schematic representation of the textual modality feature extraction model. Word embeddings are processed through dropout, Tanh activation, padding, and convolutional layers to capture multi-scale features. A residual unit is applied to ensure even questions with subtle n-gram features retain robust initial representations. Then, features are concatenated and fed into a GRU to extract sequential semantics. The inset details the GRU’s internal gating mechanisms.
@@
Figure 2 outlines our Convolutional GRU (ConvGRU) model’s architecture, highlighting the integration of convolutional n-gram feature extraction with GRU processing. Given a sequence of question words $q$, each word is first embedded into a high-dimensional space, yielding the matrix $E\in\mathbb{R}^{B\times L\times d}$, where $B$ is the batch size, $L$ is the sequence length, and $d$ is the embedding size. Following a dropout and Tanh activation, the matrix is calculated as:
@@
With this schema, our Convolutional GRU improves upon traditional methods by adeptly capturing and emphasizing pivotal textual features, thus providing a robust foundation for more accurate responses.
@@ 4.1 Dataset and Evaluation Metric
where $A$ is the human answer set including 10 answers, and $Agree(\cdot)$ is the number of human answers same with the predicted one. An answer is considered correct if at least 3 humans give the same answer to ensure accuracy while accounting for the possibility of multiple correct answers to the same question.
To address language priors and answer imbalance in the dataset, we also employ the Balanced Pairs Accuracy [26] metric. It forms complementary pairs by asking the same question about two different images requiring different answers. The model receives credit only if it answers both correctly. Mathematically, Balanced Pairs Accuracy is calculated as:
$\overline{Acc}=\frac{1}{B}\sum_{k=1}^{B}\delta(Acc_{k}^{(1)}==1\,and\,Acc_{k}^{(2)}==1)$ (16)
where $\overline{Acc}$ is the Balanced Pairs Accuracy, $B$ denotes the number of complementary pairs, $Acc_{k}^{(1)}$ and $Acc_{k}^{(2)}$ are the accuracies for two images in the $k$-th pair, computed using Equation (15). $\delta(\cdot)$ returns 1 if both accuracies are 1, and 0 otherwise.
@@ 4.2 Experimental Settings
4.2.1 Training Details
All experiments described in this study were conducted on a single RTX 3090 GPU with 24GB of memory. For the purpose of maintaining consistency and fairness across all tests, we utilized uniform parameters when evaluating a variety of Text Feature Extraction Models (under the same random seed). Specifically, the following settings were applied to each experiment:
@@ 4.2.1 Training Details
    $lr_{i+1}=lr_{i}\times 0.5^{(1/lr\_half)}$ (17)
    Here, $lr_{i}$ represents the learning rate at the current epoch, and $lr_{i+1}$ denotes the learning rate for the next epoch. The term $lr\_half$ is set to 50,000, indicating that the learning rate is halved every 50,000 epochs.
@@
4.2.2 Compared Methods
Image features are computed from R-CNN model with ResNet-101 as backbone mentioned in Section 3.3.1. Different text models are adopted, including basic one-layer Gated Recurrent Unit (GRU), Text Convolutional Neural Network (TextCNN), two-layer Long Short-Term Memory Network(2-Layer-LSTM), Bidirectional GRU (BiGRU), Bidirectional LSTM (BiLSTM), Self-Attention GRU (SA-GRU), Multi-Head Self-Attention GRU (MHA-GRU), Transformer Encoder (TE) and Transformer Encoder with GRU (TE-GRU). Specific settings are as follows: